### v4

@@ 3 Approach
Figure 2: The overall framework of SAR-JEPA. (a) In the pre-training stage, our Joint-Embedding Predictive Architecture for SAR target recognition (SAR-JEPA) uses local masked patches to predict the multi-scale SAR feature representations $f_{\rm SAR}$ of unseen patches. The MIM structure uses the Vision Transformer (ViT) in MAE and extracts deep features $f_{\rm Deep}$ of masked patches. Its DNN predictor predicts SAR features of unseen patches. The target encoder uses the gradient-by-ratio method to map SAR images from pixel to feature space, thus extracting the target shapes and avoiding speckle noise interference in SAR images. Local patches and multi-scale features are designed for multi-scale small targets in remote sensing. For downstream datasets, we fine-tune the DNN encoder with a task layer using labeled training data. DNN encoder weights are frozen or fine-tuned using different tuning settings. Other model module weights are dropped. (c) The trained model is used to predict the test result.
The proposed Joint-Embedding Predictive Architecture for SAR target recognition (SAR-JEPA) is illustrated in Figure 2. Our overall objective is as follows: given local masked patches, predict feature representations in the SAR domain, w.r.t, the model can achieve SAR image interpretation by recognizing the contextual relationship between the target parts and the surrounding area.
As shown in Figure 2, the input is the single-channel SAR magnitude image since most target datasets are magnitude pictures. We randomly select different local patches and add mask tokens according to LoMaR (Chen et al., 2022a). Vision Transformers (ViT) (Dosovitskiy et al., 2020) in MAE are used as the DNN-encoder and DNN-predictor. Each local masked patch is fed into the MIM structure to predict the target feature of mask tokens. Importantly, the special SAR coherent imaging mechanism requires predicting high-quality target features instead of noisy pixels. SAR coherent imaging results in multiplicative speckle noise in the magnitude image, and we use the gradient-by-ratio (Dellinger et al., 2014; Song et al., 2016) to suppress noise and extract the targets’ edge and shape. Moreover, we employ a multi-scale and local mask setting for various targets in the pre-training set.
@@ 3.1 Local Masked Patches
We aim to learn the contextual relationship of various targets. However, such relationships normally remain in small areas due to the large scale and small target in remote sensing imagery. Therefore, we perform masking in the local patches following LoMaR. Although LoMaR is designed to improve computational efficiency, we found this local setting is suited for small targets in remote sensing. According to the LoMaR, mask tokens and relative positional encoding (Wu et al., 2021) are added to the encoder input and MAE structure to enhance the feature representation. These are important for LoMaR to avoid feature collapse.
@@ 3.2 Target Encoder
Due to the SAR imaging mechanism (Moreira et al., 2013), the images’ amplitude values contain multiplicative speckle noise. The speckle noise is related to the sensor resolution and is caused by the coherent superposition of different scatterers within a scattering cell. Therefore, reconstructing image pixels for SAR images receives noise interference and fails to learn a high-quality feature representation. This problem of speckle noise can be solved by filtering and feature extraction. However, filtering has a conflict between noise reduction and image details, and we use feature extraction. Some studies have proposed different types of target features in computer vision, mainly divided into traditional manual features (Wei et al., 2022) and deep learning features (Assran et al., 2023). However, these methods are difficult to migrate to SAR target recognition directly, and we combine the well-established local description feature methods (Dellinger et al., 2014; Song et al., 2016; Dong et al., 2020) in the SAR domain.
Gradient computation. The differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in the SAR image, and previous studies (Touzi et al., 1988; Bovik, 1988) have shown that the computing ratio is more suitable for multiplicative noise. Here, we use the simplest ratio of average (ROA) (Touzi et al., 1988) so as not to blur image details:
@@
where $R_{i}$ denotes the average ratio at different directions and $M_{1}(i)$ and $M_{2}(i)$ are the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction and $i=3$ means the vertical direction. As shown in Figure 2, the area averages $M$ can be computed from the input image with four fixed convolution kernels [^3] We added 1e-2 to the input image to prevent zero values.. Each convolution kernel extracts the sum of the region on one side of the pixel point, which is averaged to obtain $M$. By varying the size of the convolution kernel, the multi-scale features $f_{\rm SAR}$ can be obtained.
Then, the use of logarithms can solve the vertical gradient calculation (Dellinger et al., 2014), and the horizontal and vertical gradients are defined as:
$\centering\begin{split}G_{\text{H}}&=log(R_{1})\\ G_{\text{V}}&=log(R_{3})\end{split}\@add@centering$ (2)
where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient. The gradient magnitude is $G_{\text{m}}=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}}$. We use gradient magnitude to construct multi-scale gradient features $f_{\rm SAR}$.
Multi-Scale Feature. SAR-HOG (Song et al., 2016) has discussed the kernel sizes for SAR vehicle target recognition and given the best single size for the MSTAR vehicle dataset. However, a single-size kernel cannot accommodate various targets when the dataset is extended to different targets and senors. Due to the dynamic range of contextual information for various targets in remote sensing (Li et al., 2023c), a multi-scale feature is constructed with convolutional kernels of different sizes in Figure 1.
We set the kernel size $r$ equal to 5, 7, 13, and 17, and the whole convolution kernel is a square with an odd $12r+$ 1. Though different kernel sizes, we have four gradient features $G_{\rm m1}$, $G_{\rm m2}$, $G_{\rm m3}$, and $G_{\rm m4}$, and the multi-scale feature $f_{\rm SAR}={\rm cat}(G_{\rm m1},G_{\rm m2},G_{\rm m3},G_{\rm m4})$ is a concatenat of four gradient features in the feature channel. As shown in Figure 1, small-scale convolution kernels provide a finer extraction of small target contours, while large-scale convolution kernels focus on large targets and scene edges and noise suppression.
@@ 3.3 Implementation
Given SAR images, we randomly sample several square windows at different spatial locations as local patches and mask each patch with mask tokens at a fixed percentage. Then, each local masked patch’s visible and unseen parts are fed into the MIM structure, whose encoder and predictor are applied with learnable relative positional encoding in the self-attention layer. Meanwhile, the whole SAR image is provided to the target encoder. The average values in four directions are obtained after multi-scale convolution in four directions. Then, the multi-scale SAR features are obtained after the ratio, logarithm, and magnitude operations. The loss function computes the mean squared error in the feature space between the DNN features $f_{\rm Deep}$ and SAR multi-scale features $f_{\rm SAR}$ for masked patches. In addition to the major improvement of the target encoder, there are two other minor.
@@
Decoder design. While some work (Chen et al., 2022a; Wei et al., 2022) has shown that it is feasible to use a simple linear head as a decoder or predictor for self-supervised learning in RGB images, we find that a ViT with eight-layers in the MAE is less disturbed by SAR image noise than a linear layer, allowing the encoder to extract deep semantic features of targets.
@@ 4 Experiments
This section analyzed the proposed architecture. We first described the pre-training and few-shot classification experimental setting in Sec. 4.1. Then, we discussed the effectiveness of the different modules in our architecture in Sec. 4.2. And the proposed SAR-JEPA was compared with other methods in Sec. 4.3 and in Sec. 4.4. In the end, Sec. 4.5 discussed the scaling experiment and SAR-JEPA showed improved performance with increasing unlabeled data.
@@ 4.1 Dataset and Experimental Settings
Table 1: Description of SAR datasets used for pre-training and downstream tasks. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Large SAR imagery in the detection contains more target and scene types than the annotation.
@@
MSAR 28,499 $\geq$4 $\geq$6 1 C Quad Ground and sea targets detection dataset
@@
Figure 3: Datasets for the pre-training and downstream tasks. Pre-training contain different targets, scenes, and sensors from MSAR, SAR-Ship-Dataset, SARSim, and SAMPLE. MSAR is the satellite-based dataset of ground and sea targets; SAR-Ship-Dataset is the satellite-based dataset of sea targets; SARSim is the multi-angle simulation dataset of vehicle targets; and SAMPLE is the simulated and measured dataset of vehicle targets. We use MSTAR, FUSAR-Ship, and SAR-ACD datasets to evaluate the performance for recognizing different targets. MSTAR is a fine-grained vehicle dataset; FUSAR-Ship is a sea targets dataset; SAR-ACD is a fine-grained aircraft dataset.
@@
We examine self-supervised learning performance with the following procedure. First, we perform the self-supervised learning on the pre-training dataset without label information. Then, we fine-tune the pretrained model on few-shot SAR classification datasets.
@@
Pre-training. Pre-training datasets are designed to ensure rich information about targets, scenes, and sensors. Therefore, we chose four datasets to build a pre-training dataset for SSL. It covers popular targets (vehicles, ships, aircraft, etc.), complex backgrounds (ground and ocean), and various sensors (satellite-borne, airplane-borne, and simulation) in Table 1 and Figure  3.
@@
MSAR (Xia et al., 2022; Chen et al., 2022b) is a multi-class target detection dataset based on the Chinese HISEA-1 satellite in large-scale scenes. MSAR comprises 28,449 image slices with quad polarization. Scenes covered include airports, harbors, nearshore, islands, distant seas, urban areas, etc. The labeled target categories include aircraft, oil tanks, bridges, and ships. This dataset can provide target samples in complex ground and sea scenes.
@@
SAR-Ship (Wang et al., 2019) dataset is a ship detection dataset in complex scenes based on Chinese Gaofen-3 and European Sentinel-1 satellites. The public version of this dataset contains 39,729 images from two satellites in different imaging modes and resolutions. The dataset provides ship targets of various sizes in complex ocean scenes such as nearshore, distant seas, harbors, and islands.
@@
SARSim (Malmgren-Hansen et al., 2017; Kusk et al., 2016) is a fine-grained vehicle dataset created by Terma A/S, Denmark. The simulation system used for this dataset can generate X-band SAR images with resolutions ranging from 0.1m to 0.3m from CAD models. SARSim provides 21,168 vehicle samples in 7 categories (truck, car, motorbike, bus, tank, bulldozer, and pickup) and 3 scenes (grass, roads, and a mean of the two) with 7 imaging depression angles. This dataset provides fine-grained differences between targets.
@@
SAMPLE (Lewis et al., 2019) is a fine-grained synthetic and measured paired vehicle dataset released by the Air Force Research Laboratory, USA. This dataset is simulated in X-band and 0.3m resolution. The public version of this dataset provides 5,380 images of ten categories of military vehicle targets at partial imaging angles, and five categories overlap with the public version of MSTAR. As shown in Figure 3, it can be noticed that the ground clutter in SAMPLE is more significant and realistic than in SARSim.
@@
Downstream tasks. We tested the pre-trained model on three target recognition datasets to evaluate the ability for fine-grained semantic classification in the case of small samples with fine-tuning and linear probing. According to MAE, fine-tuning is tuning all model parameters, and linear probing is tuning a linear layer with batch normalization while freezing other parameters.
@@
MSTAR (Air Force Research Laboratory, ) is the most commonly used SAR target classification dataset released by the Defense Advanced Research Projects Agency, USA, which contains ten categories of vehicles with different angles, variants, and other imaging conditions. It has many experimental setting variants, while we refer to the (Li et al., 2023b) to adopt the most commonly used ten-class classification settings to evaluate the fine-grained classification performance for SAR vehicles.
@@
FUSAR-Ship (Sun et al., 2022) contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island. Based on the experimental setting of (Wang et al., 2022a), we have ten types of ocean target slices from origin images, such as four fine-grained ships, bridges, and ocean scene slices. We use this dataset to evaluate the fine-grained classification performance for ocene targets.
@@
SAR-ACD (Sun et al., 2022) contains six types of aircraft based on the Gaofen-3 satellite in three civil airports, including Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport. Since the released dataset does not separate the training and test data, we randomly select small samples as the training set and others as the test set. We use this dataset to evaluate fine-grained aircraft target classification. Fine-grained recognition of aircraft targets is a more challenging task due to the smooth surface of the aircraft, resulting in insignificant features.
@@
Experimental settings. TTable 2 gives the default setting. The pre-training is applied on 4 NVIDIA V100 GPUs with 200 epochs and 300 batch sizes, and the backbone model is ViT-Base (ViT-B). Other specific hyperparameters of each method use the default settings from their papers and codes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness, and all methods use the same data augmentation. Since our goal is not to get better performance by tuning downstream hyperparameters, all models use the same training settings in downstream classification tasks. The few-shot learning setting is based on the Dassl toolbox (Zhou et al., 2023a) and is averaged over 10 random experiments.
Table 2: Experimental setting.
Pre-training setting

### v5

@@ 3 Approach
Figure 2: Overall framework of SAR-JEPA. (a) In the pretraining stage, joint-embedding predictive architecture for SAR automatic target recognition (SAR-JEPA) uses local masked patches to predict the multi-scale SAR feature representations $f_{\rm SAR}$ of unseen patches. The MIM structure uses the Vision Transformer (ViT) in MAE to extract deep features $f_{\rm Deep}$ of masked patches. Its DNN predictor predicts the SAR features of unseen patches. The target encoder uses the GR method to map SAR images from pixel to feature space, thus extracting the target shapes and avoiding speckle noise interference in SAR images. Local patches and multi-scale features are designed for multi-scale small targets in remote sensing. For downstream datasets, we fine-tune the DNN encoder with a task layer using labeled training data. DNN encoder weights are frozen or fine-tuned using various tuning settings. The weights of the other model modules are removed. (c) The trained model is used to predict the test result.
The proposed SAR-JEPA is illustrated in Figure 2. Our general objective is as follows: given local masked patches, predicting feature representations in the SAR domain with respect to the model can interpret the SAR image by recognizing the contextual relationship between the target parts and the surrounding area.
As shown in Figure 2, the input is the single-channel SAR magnitude image because most target datasets are magnitude pictures. We randomly select different local patches and add mask tokens based on LoMaR (Chen et al., 2022a). ViTs (Dosovitskiy et al., 2020) in MAE are used as the DNN-encoder and DNN-predictor. Each local masked patch is fed into the MIM structure to predict the target feature of the mask tokens. Importantly, the special SAR coherent imaging mechanism requires predicting high-quality target features instead of noisy pixels. SAR coherent imaging results in multiplicative speckle noise in the magnitude image, and we use GR (Dellinger et al., 2014; Song et al., 2016) to suppress noise and extract the target shapes. Moreover, we employ a multi-scale and local mask setting for various targets in the pretraining set.
@@ 3.1 Local Masked Patches
We aim to learn the contextual relationship between various objects. However, such relationships usually remain in small areas because of the large scale and small target in remote sensing imagery. Therefore, we perform masking on the local patches following LoMaR. Although LoMaR is designed to improve computational efficiency, we found that this local setting is suitable for small targets in remote sensing. According to LoMaR, mask tokens and relative positional encoding (Wu et al., 2021) are added to the encoder input and MAE structure to improve the feature representation. These are important for LoMaR to avoid feature collapse.
@@ 3.2 Target Encoder
Owing to the SAR imaging mechanism (Moreira et al., 2013), the image amplitude values contain multiplicative speckle noise. Speckle noise is related to the sensor resolution and is caused by the coherent superposition of different scatterers within a scattering cell. Therefore, reconstructing image pixels for SAR images suffers from noise interference and fails to learn a high-quality feature representation. This speckle noise problem can be solved using filtering and feature extraction. However, filtering has a conflict between noise reduction and image details, and we use feature extraction. Some studies have proposed various types of target feature in computer vision, which are mainly divided into traditional manual features (Wei et al., 2022) and deep learning features (Assran et al., 2023). However, these methods cannot easily migrate directly to SAR target recognition, and we combine well-established local description feature methods (Dellinger et al., 2014; Song et al., 2016; Dong et al., 2020) in the SAR domain.
Gradient computation. The differential gradient is not a constant false alarm rate operator because of the multiplicative speckle noise in the SAR image. Previous studies (Touzi et al., 1988; Bovik, 1988) have shown that the computing ratio is more suitable for multiplicative noise. Here, we use the simplest ratio of average (ROA) (Touzi et al., 1988) to avoid blurred image details:
@@
where $R_{i}$ denotes the average ratio at different directions and $M_{1}(i)$ and $M_{2}(i)$ denote the area averages on opposite sides of the current pixel along direction $i$. $i=1$ is the horizontal direction, and $i=3$ is the vertical direction. As shown in Figure 2, the area averages $M$ can be computed from the input image using four fixed convolution kernels [^2] We added 1e-2 to the input image to prevent zero values.. Each convolution kernel extracts the sum of the region on one side of the pixel point, which is averaged to obtain $M$. By varying the sizes of the convolution kernel, multi-scale features $f_{\rm SAR}$ can be obtained.
Then, the use of logarithms can solve the vertical gradient calculation (Dellinger et al., 2014), and the horizontal and vertical gradients are defined as follows:
$\centering\begin{split}G_{\text{H}}&=log(R_{1}),\\ G_{\text{V}}&=log(R_{3}),\end{split}\@add@centering$ (2)
where $G_{\text{H}}$ denotes the horizontal gradient and $G_{\text{V}}$ denotes the vertical gradient. The magnitude of the gradient is $G_{\text{m}}=\sqrt{G_{\text{H}}^{2}+G_{\text{V}}^{2}}$. We use gradient magnitude to construct multi-scale gradient features $f_{\rm SAR}$.
Multi-Scale Feature. SAR-HOG (Song et al., 2016) discusses the kernel sizes for the MSTAR vehicle and provides the best single size for the MSTAR vehicle dataset. However, a single-sized kernel cannot accommodate various targets when the dataset is extended to different targets and sensors. Owing to the dynamic range of contextual information for various targets in remote sensing (Li et al., 2023c), a multi-scale feature is constructed with convolutional kernels of various sizes, as depicted in Figure 1.
We set the kernel size $r$ equal to 5, 9, 13, and 17, and the whole convolution kernel is a square with an odd number $12r+$ 1. Using various kernel sizes, we have four gradient features, $G_{\rm m1}$, $G_{\rm m2}$, $G_{\rm m3}$, and $G_{\rm m4}$, and the multi-scale feature $f_{\rm SAR}={\rm cat}(G_{\rm m1},G_{\rm m2},G_{\rm m3},G_{\rm m4})$ is a concatenate of four features in the feature channel. As shown in Figure 1, small-scale convolution kernels provide a finer extraction of small target contours, while large-scale convolution kernels focus on large targets, scene edges, and noise suppression.
@@ 3.3 Implementation
Given SAR images, we randomly sample several square windows at different spatial locations as local patches and mask each patch with mask tokens at a fixed percentage. Then, the visible and unseen parts of each local masked patch are fed into the MIM structure, whose encoder and predictor are applied with learnable relative positional encoding in the self-attention layer. Moreover, the entire SAR image is provided to the target encoder. The average values in four directions are obtained after multi-scale convolution in four directions. Then, the multi-scale SAR features are obtained after performing the ratio, logarithm, and magnitude operations. The loss function computes the mean squared error in the feature space between the DNN features $f_{\rm Deep}$ and SAR multi-scale features $f_{\rm SAR}$ for masked patches. In addition to the major improvement in the target encoder, there are two other minor ones.
@@
Decoder design. Although some studies (Chen et al., 2022a; Wei et al., 2022) have shown that it is feasible to use a linear head as a decoder for SSL in RGB images, we find that a ViT with eight layers in the MAE is less disturbed by SAR image noise than a linear layer, thus allowing the encoder to extract deep semantic features of targets.
@@ 4 Experiments
This section analyzes the proposed architecture. We first describe the experimental setting for pretraining and few-shot classification in Subsection 4.1. Then, we discuss the effectiveness of the different modules in our architecture in Subsection 4.2. The proposed SAR-JEPA was compared with other methods in Subsections 4.3 and 4.4. In the end, Subsection 4.5 discusses the scaling experiment, and SAR-JEPA shows improved performance as the amount of unlabeled data increases.
@@ 4.1 Dataset and Experimental Settings
Table 1: Description of SAR datasets used for pretraining and downstream tasks. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Large SAR imagery in the detection contains more target and scene types than the annotation.
@@
MSAR 28,499 $\geq$4 $\geq$6 1 C Quad Ground and sea target detection dataset
@@
Figure 3: Datasets for the pretraining and downstream tasks. Pretraining contains various targets, scenes, and sensors from MSAR, SAR-Ship-Dataset, SARSim, and SAMPLE. MSAR is the satellite-based dataset of ground and sea targets; SAR-Ship-Dataset is the satellite-based dataset of sea targets; SARSim is the multi-angle simulation dataset of vehicle targets; and SAMPLE is the simulated and measured dataset of vehicle targets. We use MSTAR, FUSAR-Ship, and SAR-ACD datasets to evaluate the performance in recognizing different targets. MSTAR is a fine-grained vehicle dataset; FUSAR-Ship is a sea target dataset; and SAR-ACD is a fine-grained aircraft dataset.
@@
We examine SSL performance using the following procedure. First, we perform SSL on the pretraining dataset without label information. Then, we fine-tune the pretrained model on few-shot SAR classification datasets.
@@
Pretraining. Pretraining datasets are designed to provide rich information about targets, scenes, and sensors. Therefore, we chose four datasets to build a pretraining dataset for SSL. It covers popular targets (vehicles, ships, aircraft), complex backgrounds (ground and ocean), and various sensors (satellite-borne, airplane-borne, and simulation), as illustrated in Table 1 and Figure  3.
@@
MSAR (Xia et al., 2022; Chen et al., 2022b) is a multi-class target detection dataset based on the Chinese HISEA-1 satellite in large-scale scenes. MSAR comprises 28,449 image slices with quad polarization. The scenes covered include airports, harbors, nearshore, islands, distant seas, and urban areas. Labeled target categories include aircraft, oil tanks, bridges, and ships. This dataset provides target samples in complex ground and sea scenes.
@@
SAR-Ship (Wang et al., 2019) dataset is a ship detection dataset in complex scenes based on Chinese Gaofen-3 and European Sentinel-1 satellites. The public version of this dataset contains 39,729 images from two satellites with various imaging modes and resolutions. The dataset provides ship targets of various sizes in complex ocean scenes, such as nearshore, distant seas, harbors, and islands.
@@
SARSim (Malmgren-Hansen et al., 2017; Kusk et al., 2016) is a fine-grained vehicle dataset created by Terma A/S, Denmark. The simulation system used for this dataset can generate X-band SAR images with resolutions ranging from 0.1 m to 0.3 m from computer-aided design (CAD) models. SARSim provides 21,168 vehicle samples in 7 categories (truck, car, motorbike, bus, tank, bulldozer, and pickup) and 3 scenes (grass, roads, and a mean of the two) with 7 imaging depression angles. This dataset provides fine-grained differences between targets.
@@
SAMPLE (Lewis et al., 2019) is a fine-grained synthetic and measured paired vehicle dataset released by the Air Force Research Laboratory, USA. This dataset is simulated in X-band and 0.3 m resolution. The public version of this dataset provides 5,380 images of 10 categories of military vehicle targets at partial imaging angles, and 5 categories overlap with the public version of MSTAR. As shown in Figure 3, it can be observed that the ground clutter in SAMPLE is more significant and realistic than in SARSim.
@@
Downstream tasks. We tested the pretrained model on three target recognition datasets to evaluate its ability for fine-grained semantic classification in the case of small samples with fine-tuning and linear probing. Based on MAE, fine-tuning is tuning all model parameters, and linear probing is tuning a linear layer with batch normalization while freezing the other parameters.
@@
MSTAR (Air Force Research Laboratory, ) is the most commonly used SAR target classification dataset released by the Defense Advanced Research Projects Agency, USA, and it contains ten categories of vehicles with different angles, variants, and other imaging conditions. It has many variants of experimental settings, and we refer to (Li et al., 2023b) to adopt the most widely used ten-class classification settings to evaluate the fine-grained classification performance of SAR vehicles.
@@
FUSAR-Ship (Sun et al., 2022) contains 15 primary ship categories and many nonship targets based on the Gaofen-3 satellite in scenes, such as sea, land, coast, river, and island. Based on the experimental setting (Wang et al., 2022a), we have ten types of ocean target slices from the original images, such as four fine-grained ships, bridges, and ocean scene slices. We use this dataset to evaluate the fine-grained classification performance of ocene targets.
@@
SAR-ACD (Sun et al., 2022) contains six types of aircraft based on the Gaofen-3 satellite in three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport. Because the released dataset does not separate the training and test data, we randomly select small samples as the training set and the others as the test set. We use this dataset to evaluate fine-grained aircraft target classification. The fine-grained recognition of aircraft targets is a more challenging task owing to the smooth surface of the aircraft, resulting in insignificant features.
@@
Experimental settings. Table 2 presents the default setting. The pretraining is applied on 4 NVIDIA V100 GPUs with 200 epochs and 300 batch sizes, and the backbone model is ViT-Base (ViT-B). Other specific hyperparameters of each method use the default settings of their papers and codes. Compared to the training settings of MAE, we add ColorJitter [^3] ColorJitter is a Pytorch data augmentation method for adjusting an image’s brightness, contrast, saturation, and hue. Considering the strong and weak scattering points in SAR magnitude images, adjusting contrast can effectively increase data diversity (Geng et al., 2023). Therefore, we perform the contrast adjustment through a one-line Pytorch code “transforms.ColorJitter (contrast=0.5)”. to increase the data richness, and all methods use the same data augmentation. Because our goal is not to obtain better performance by tuning downstream hyperparameters, all the models use the same training settings in downstream classification tasks. The few-shot learning setting is based on the Dassl toolbox (Zhou et al., 2023a) and is averaged over 10 random experiments.
Table 2: Experimental setting
Pretraining setting